# Evaluating Generative Ad Hoc Information Retrieval

### v2

@@ Abstract.
Recent advances in large language models have enabled the development of viable generative retrieval systems. Instead of a traditional document ranking, many generative retrieval systems directly return a grounded generated text as an answer to an information need expressed as a query or question. Quantifying the utility of the textual responses is essential for appropriately evaluating such generative ad hoc retrieval. Yet, the established evaluation methodology for ranking-based retrieval is not suited for reliable, repeatable, and reproducible evaluation of generated answers. In this paper, we survey the relevant literature from the fields of information retrieval and natural language processing, we identify search tasks and system architectures in generative retrieval, we develop a corresponding user model, and we study its operationalization. Our analysis provides a foundation and new insights for the evaluation of generative retrieval systems, focusing on ad hoc retrieval.
@@ 1. Introduction
The development of large language models (LLMs) has prompted search engine and AI companies to innovate the way search results are presented: using LLMs to directly generate a textual answer that satisfies an information need. While LLMs can generate unreliable information (Alkaissi and McFarlane, 2023; Ji et al., 2023; Koopman and Zuccon, 2023), conditioning their inference on relevant documents has emerged as a potential technique to ground their generated statements (Lewis et al., 2020; Mialon et al., 2023). This may relieve users of the (cognitive) effort of acquiring the needed information from individual search results themselves, which affords a change in the design of a search engine results page (SERP; Figure 1): instead of the proverbial list of “ten blue links” (list SERP, left), a generated text with references is shown (text SERP, right). The first public prototypes of this kind were You.com’s You Chat and Neeva AI, closely followed by Microsoft’s Bing Copilot, Google’s Gemini, Perplexity.ai, Baidu’s Ernie,111See https://chat.you.com; Neeva has shutdown; https://chat.bing.com; https://gemini.google.com; https://perplexity.ai; https://yiyan.baidu.com. and other research prototypes (Koopman et al., 2023; Zhang and Pradeep, 2023).
@@
Far ahead of this development, Sakai et al. (2011) raised an important question: How can search engines that use text SERPs be evaluated? Evaluating text SERPs is not straightforward, since the modern theory and practice of evaluation in information retrieval is premised on the assumption that search results are presented as list SERPs.222Extensive research on search interfaces has included many alternatives, search features, interaction designs, and result visualizations (Hearst, 2009; Wilson, 2011; Liu et al., 2021). Nonetheless, with the growth of the Web, the list SERP design became a de facto standard for web search.
According to list SERP user models, a ranked list of results triggers a certain user behavior like reading the results in order until the information need is satisfied or the search is abandoned. In decades of research, a comprehensive theoretical framework of reliable and validated evaluation methods has been built to assess the quality of document rankings with respect to information needs. Replacing ranked results by a generated text undermines this foundation.
In this paper, we focus on questions related to transferring established list SERP evaluation methodology to text SERPs. Our approach is theory-driven and based on a systematic analysis of relevant literature from information retrieval (IR) and related fields. Our contributions relate to the system, user, and evaluation perspectives. Starting with a definition of what generative ad hoc retrieval is, we distinguish two fundamental system models for generative retrieval and contextualize them in Broder’s ((2002)) taxonomy of search tasks (Section 2). We then devise a user model for text SERPs, grounded in related behavioral studies (Section 3). Finally, we revisit IR evaluation methodologies to develop a foundation for text SERP effectiveness measures and for the reliable evaluation of generative ad hoc retrieval (Section 4).
2. Generative Retrieval Tasks
@@ 2.1. Generative Ad Hoc Retrieval
The tasks of language generation and of retrieval appear to be distinct. However, retrieval systems and generative language models are both created using large collections of documents (see Figure 2 top), and the usefulness of both depends on tuning them with user needs, expressed as queries or prompts . The users of an IR system want to retrieve the most relevant documents for a query, and the users of a generative language model want to generate the most helpful text for a prompt. From an IR perspective, the most salient difference between retrieval and generation is as follows: a retrieval model induces a ranking on a finite document collection , while a generative language model induces a ranking on the infinite set of all possible texts . In practice, retrieval models are used to return the top- ranked documents, and generative language models to return just one of the many possible relevant texts from . Generative models have therefore recently been framed as infinite indexes (Deckers et al., 2023).
@@
As a retrieval model can only return existing documents, the information available in determines the degree to which a user’s information need can be satisfied. Still, the user has to examine the returned documents for the desired information. A generative language model instead attempts to alleviate the effort of examining documents by returning a tailored response that compiles all desired information. Yet, the factual accuracy of current generative language models is often lacking and prone to confabulations or hallucinations (Zhao et al., 2023; Koopman and Zuccon, 2023; Ji et al., 2023; Alkaissi and McFarlane, 2023) (i.e., there is only a very small subset of accurate texts among all possible texts ).333For cases like counterfactual information needs (e.g., “What if Columbus didn’t settle America?” (Kiciman et al., 2023)), strong confabulation capabilities could be explicitly desirable, though.
@@
Generative ad hoc retrieval can be described as combining retrieval and generation in ad hoc scenarios (one query, one result) in a way that the used systems’ respective advantages complement each other. The two fundamental combination approaches are (Figure 2, bottom): retrieving relevant documents from and generating an answer from them, or generating a response and “verifying” its statements by retrieving supporting documents from .
2.2. Operationalization Paradigms
@@
System for generative ad hoc retrieval require two components (Gao et al., 2023): (1) retrieval, to gather existing documents from a collection for a query, and (2) generation, to generate a text for a prompt. Both components can be combined in two ways (Gao et al., 2023): retrieval-then-generation and generation-then-retrieval, which form the two paradigms of operationalizing generative ad hoc retrieval (Figure 2, bottom). In a retrieval-then-generation approach, a language model is conditioned with retrieved source material, for instance, by adding evidence to its input prompt (Izacard and Grave, 2021; Khot et al., 2023; Lazaridou et al., 2022; Shi et al., 2023), attending to retrieved sources during inference (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022), chaining ideas (Jiang et al., 2022), or iterative self-attention starting from sources (Zhang et al., 2021b). In a generation-then-retrieval approach, the retrieval model is used to find sources for generated text passages. Though this approach has received little attention (AutoGPT Contributors, 2023), it resembles retrieving references for individual generated statements, similar to claim verification (Wadden et al., 2020).
@@
With increasing inference speeds of generative language models, arbitrarily ordered combinations of the retrieval and the generation step are possible, leading to multi-step generative ad hoc retrieval. The simplest form might be iterative cyclical patterns like generating a text passage that is used as a query to retrieve relevant sources, which in turn serve as context for the next iteration, etc. Applications are the continuous generation of text (Ram et al., 2023; Jiang et al., 2023; Semnani et al., 2023), retrieving sources in multiple steps (Patwardhan and Yang, 2023), or the refinement of a text through iterative inference (Khattab et al., 2022; AutoGPT Contributors, 2023).
@@
In this paper, we focus on the evaluation of the text SERP output of (multi-step) generative ad hoc retrieval, but do not consider evaluating any step individually.
@@ 2.3. Generative (Ad Hoc) Search Tasks
Ad hoc retrieval has a long history with a large body of respective research (Manning et al., 2008). Its goal is to output a single result (ranking) for a single-query input (i.e., the information need must be satisfied without knowing any previous queries or interactions). In 2002, Broder suggested a now well-known taxonomy of search tasks (Broder, 2002) and related them to three generations of web search systems (see Table 1). Each generation utilizes a new source of information in addition to those of its predecessors to meet new user intents. The first generation supports informational tasks, relying only on the information found within single documents to support a user’s intent to acquire (parts of) that information. The second generation additionally exploits document relations, supporting users to reach a specific site, document, or the most authoritative one among many alternatives (i.e., navigational tasks). The third generation blends results from different vertical systems and multimodal results into a single SERP to support a user in performing transactional tasks.
Generative retrieval systems can be seen as a new, 4th generation of web search systems. They enable the synthesis of new “documents” relevant to a user’s information need. Given a sufficiently complex need that cannot be satisfied by returning the URL of a web page, generative retrieval systems are primarily used to synthesize a comprehensive overview of the information required, condensed into a long-form answer. Prior search system generations delegated the condensation of information from multiple sources to the users who then often needed to study several documents from a list SERP to satisfy their needs. Generative models promise to reduce this extra work and cognitive load, so that users need only read one generated text.444Already in (2011), Sakai et al. (2011) have previously proposed to automatically identify relevant information nuggets in retrieved documents and to only present the nuggets in a list, but did not consider the aspect of condensing them. Additionally, the “synthesizing” nature of generative retrieval systems can conceivably be harnessed to generate new pieces of information not contained in the retrieved sources, rendering the generative model itself a source of information.
@@
While the search tasks addressed by generative retrieval systems are in many cases informational in nature, we argue to consolidate them as synthetic search tasks. Typical examples are opinion needs (“Should society invest in renewable energy?”) or decision-making needs (“Should I get life insurance?”). Such needs are not fully supported by the first three system generations: in contrast to (1) informational tasks, the information is likely spread across multiple documents; in contrast to (2) navigational tasks, no single page is premeditated to be reached by the user; and in contrast to (3) transactional tasks, the goal of condensing the information should be addressed on the system side. As if foreseeing synthetic retrieval, Broder explicitly constrained informational queries and first-generation systems to static content: “The purpose of such [informational] queries is to find information assumed to be available on the Web in a static form. No further interaction is predicted, except reading. By static form we mean that the target document is not created in response to the user query.” (Broder, 2002, page 5)
The fourth generation of web search engines supports synthetic search tasks and enables users to access a single, comprehensive generated answer (document) that satisfies a complex need with an in-depth analysis covering multiple perspectives. Although the Web may offer a static (set of) document(s) to satisfy such a need, an ideal generative retrieval system retrieves them, synthesizes missing information, presents a coherent answer, and grounds it in the retrieved sources, i.e., dynamically addresses the query.
@@ 2.4. A Taxonomy of Generative Retrieval
‘Generative retrieval’ or ’generative IR’ are umbrella terms for a diversity of approaches that use generative models to solve retrieval tasks.555See also the recent SIGIR workshop on generative IR (Bénédict et al., 2023). Following Arora et al. (2023), Figure 3 categorizes these approaches into generation-augmented retrieval (GAR) and retrieval-augmented generation (RAG). Notably, GAR approaches create traditional list SERPs, while RAG approaches generate text SERPs.
In GAR approaches, generative models are used to enhance the traditional search architecture, which can be done at indexing time and at query time. At indexing time, generative models can be used for augmenting documents (Nogueira et al., 2019; Gospodinov et al., 2023; MacAvaney et al., 2020; Formal et al., 2021; Zhuang and Zuccon, 2021) with confabulated or hallucinated content, or for replacing the standard indexing process with what are commonly termed ‘differentiable indices’ by, for instance, generating document identifiers like page titles (De Cao et al., 2021; Thorne, 2022; Chen et al., 2022), URLs (Ziems et al., 2023), or (structured) string identifiers (Zhou et al., 2023; Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022). At query time, generative models can be used for augmenting queries (MacAvaney et al., 2021; Alaofi et al., 2023), or for modeling relevance by, for instance, generating parts of existing documents from the query and retrieval by string matching (Bevilacqua et al., 2022), predicting a (re-)ranking directly (Sun et al., 2023), or using special tokens as relevance signal (Nogueira et al., 2020; Ma et al., 2023; Qin et al., 2023; Zhuang et al., 2023).
In RAG approaches—the focus of our paper—, generative models are augmented with retrieval capabilities; either internally as what we term ‘attention-level RAG’, where the context attended to during generation is retrieved concurrently (Guu et al., 2020; Jiang et al., 2022; Borgeaud et al., 2022), or externally as ‘prompt-level RAG’, where the retrieved context is inserted into the prompt. Orthogonally, one can distinguish the RAG variants retrieval-then-generation and generation-then-retrieval.
Beyond GAR and RAG, generative models can also be used to directly generate a response without relying on retrieved information (Robinson and Wingate, 2023), i.e., as infinite indexes (Deckers et al., 2023). This may involve generating multiple candidates and selecting the best one or regenerating a new response conditioned on the previous ones (Yu et al., 2023). Moreover, an answer to a generative ad hoc request can also be the first turn of a conversational search (Salton, 1969; Radlinski and Craswell, 2017; Culpepper et al., 2018), where generative models have led to new tools (Zhang et al., 2021a; Miller et al., 2017) and dialog options (Zamani et al., 2020).
@@ 3. A User Model for Generative IR
Any IR system should align with user expectations, thus, evaluation needs to be grounded in a user model. Yet, existing user models that have been derived to facilitate evaluation in IR are based on the assumptions of list SERPs. After preliminary considerations to derive a text SERP user model (Section 3.1), we first consider the general search process of a user and explore how it relates to generative approaches (Section 3.2). Then, we follow the evaluation methodology proposed by Agosti et al. (2014): first, we define evaluation objectives (Section 3.3) and then we devise a user model that corresponds to these objectives (Section 3.4). This makes it possible to later derive metrics that operationalize the user model.
The general structure is reflected in Figure 4. We base our proposed evaluation methodology on the ad hoc information search process (Vakkari, 2016) as seen from the users’ perspective (top of the figure), and formulate evaluation objectives that correspond to each component (bottom of the figure), which take into account the evaluation setting from which a user model can be induced. Traditional IR can assist the user only during Selection with a list SERP. Meanwhile, generative retrieval encompasses all three steps of Selection, Interaction, and Synthesis, to support the synthetic search task and respond with a text SERP. An evaluation of a generative retrieval system should therefore focus on these steps, which are mirrored in the Retrieval, Grounding, and Presentation evaluation objectives.
The figure consist of two aligned rows. The upper row illustrates the user information search process, with the four components of Formulation, Selection, Interaction, and Synthesis. The lower row illustrates generative IR evaluation objectives, with the four components Prompting, Retrieval, Grounding, and Presentation. Three pairs of components are related: Selection and Retrieval, Interaction and Grounding, as well as Synthesis and Presentation. These three pairs represent the steps supported by generative IR systems.

### v3

@@ Abstract.
Recent advances in large language models have enabled the development of viable generative retrieval systems. Instead of a traditional document ranking, generative retrieval systems often directly return a grounded generated text as a response to a query. Quantifying the utility of the textual responses is essential for appropriately evaluating such generative ad hoc retrieval. Yet, the established evaluation methodology for ranking-based ad hoc retrieval is not suited for the reliable and reproducible evaluation of generated responses. To lay a foundation for developing new evaluation methods for generative retrieval systems, we survey the relevant literature from the fields of information retrieval and natural language processing, identify search tasks and system architectures in generative retrieval, develop a new user model, and study its operationalization.
@@ 1. Introduction
The development of large language models (LLMs) has prompted search engines to innovate the way results are presented: using LLMs to directly generate a textual response from a query’s results. While LLMs can generate unreliable information (Alkaissi and McFarlane, 2023; Ji et al., 2023; Koopman and Zuccon, 2023), conditioning their inference o n relevant search results has emerged as a potential technique to ground generated statements (Lewis et al., 2020; Mialon et al., 2023). As textual answers can relieve users of the (cognitive) effort of collecting the needed information from individual search results themselves, the design of some search engine’s results pages (SERPs) has changed (Figure 1): instead of the proverbial list of “ten blue links” (list SERP, left), a generated text with references is shown (text SERP, right). The first public prototypes of this kind were You.com’s You Chat and Neeva AI, closely followed by Microsoft’s Bing Copilot, Google’s Gemini, Perplexity.ai, Baidu’s Ernie,111See https://chat.you.com; Neeva has shut down; https://chat.bing.com; https://gemini.google.com; https://perplexity.ai; https://yiyan.baidu.com. and other research prototypes (Koopman et al., 2023; Zhang and Pradeep, 2023).
@@
Far ahead of this development, already in (2011) Sakai et al. (2011) raised an important question: how can text SERP-based search engines be evaluated? An answer was and is not that straightforward, since the modern theory and practice of retrieval evaluation is premised on the assumption that search results are presented as list SERPs.222Even though research on search interfaces has suggested and studied many interaction designs and variants of result presentation (Hearst, 2009; Wilson, 2011; Liu et al., 2021), with the growth of the Web, the list SERP design became a de facto standard for web search.
According to list SERP user models, a ranked list of results triggers a certain user behavior like reading the results in order until the information need is satisfied or the search is abandoned. In decades of research, a comprehensive theoretical framework of reliable and validated evaluation methods has been built to assess the quality of result rankings with respect to information needs. Replacing ranked results by a generated text undermines this foundation.
In this paper, we focus on questions related to transferring established list SERP evaluation methodology to text SERPs. Our approach is theory-driven and based on a systematic analysis of relevant literature from information retrieval (IR) and related fields. Our contributions relate to the system, user, and evaluation perspectives. Starting with a definition of what generative ad hoc retrieval is, we distinguish two fundamental system models for generative retrieval and contextualize them in Broder’s (Broder, 2002) taxonomy of search tasks (Section 2). We then devise a user model for text SERPs, grounded in related behavioral studies (Section 3). Finally, we revisit IR evaluation methodologies to develop a foundation for text SERP effectiveness measures and for the reliable evaluation of generative ad hoc retrieval (Section 4).
2. Generative Retrieval
@@ 2.1. Generative Ad Hoc Retrieval
Ad hoc retrieval refers to scenarios where a user submits a single query and expects the underlying information need to be satisfied by a single result set (i.e., the information need must be satisfied without knowing any previous queries or interactions). At first glance, this ad hoc retrieval task and the task of language generation seem to be quite different. However, retrieval systems and generative language models are both built using document collections (see Figure 2, top), and the usefulness of both depends on tuning them with user needs, expressed as queries or prompts . Users of a retrieval system want to retrieve the most relevant documents for a query, and users of a generative language model want it to generate the most helpful text for a prompt. From an IR perspective, the most salient difference is that a retrieval model induces a ranking on a finite document collection , while a generative language model induces a ranking on the infinite set of all possible texts ; generative models were thus recently also framed as infinite indexes (Deckers et al., 2023). In practice, though, retrieval models only return the top- results, and generative language models only return one of the possibly many relevant texts from .
@@
As a retrieval model can only return existing documents, the information available in the underlying collection determines the degree to which a user’s information need can be satisfied. Still, the user has to examine the returned documents for the desired information. A generative language model instead attempts to alleviate the effort of examining documents by returning a tailored response that integrates all desired information. Yet, the factual accuracy of current generative language models is often prone to confabulations or hallucinations (Zhao et al., 2023; Koopman and Zuccon, 2023; Ji et al., 2023; Alkaissi and McFarlane, 2023) (i.e., there is only a very small subset of accurate texts among all possible texts ).333For counterfactual information needs (e.g., What if Columbus didn’t discover America? (Kiciman et al., 2023)), strong confabulation capabilities could be explicitly desirable, though.
@@
The term ‘generative ad hoc retrieval’ refers to approaches that combine the advantages of retrieval and generation in ad hoc scenarios (one query, one result) by retrieving relevant documents from and generating an answer from them, or by generating a response and “verifying” its statements by retrieving supporting documents from (Figure 2, bottom left resp. right).
2.2. Two Operationalization Paradigms
@@
Systems for generative ad hoc retrieval require a retrieval component to gather existing documents from a collection for a query, and a generation component to generate a text for a prompt. These components can be combined following two different paradigms (Gao et al., 2023): retrieval-then-generation or generation-then-retrieval (Figure 2, bottom). In a retrieval-then-generation approach, a language model is conditioned with retrieved source material, for instance, by adding evidence to its input prompt (Izacard and Grave, 2021; Khot et al., 2023; Lazaridou et al., 2022; Shi et al., 2023), by attending to retrieved sources during inference (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022), by chaining ideas (Jiang et al., 2022), or by iterative self-attention (Zhang et al., 2021b). In a generation-then-retrieval approach, the retrieval model is used to find sources for generated text passages. Though this idea has received less attention so far (AutoGPT Contributors, 2023), it resembles retrieving references for individual generated statements, similar to claim verification (Wadden et al., 2020).
@@
With increasing inference speeds of generative language models, arbitrarily ordered combinations of multiple retrieval and generation steps are possible, leading to multi-step generative ad hoc retrieval. The simplest form might be iterative cycles like generating a text passage that is used as a query to retrieve relevant sources, which in turn serve as context for the next generation, etc. Applications are the continuous generation of text (Ram et al., 2023; Jiang et al., 2023; Semnani et al., 2023), retrieving sources in multiple steps (Patwardhan and Yang, 2023), or the refinement of a text through iterative inference (Khattab et al., 2022; AutoGPT Contributors, 2023).
@@
In this paper, we focus on the evaluation of the text SERP output of (possibly multi-step) generative ad hoc retrieval, but we do not consider evaluating any step individually.
@@ 2.3. Generative (Ad Hoc) Search Tasks
In 2002, Broder suggested a now well-known taxonomy of search tasks (Broder, 2002) and related them to three generations of web search systems (see Table 1). Each generation utilizes a new source of information in addition to those of its predecessors to meet new user intents. First-generation systems support informational tasks, relying only on the information found within some single document to support a user’s intent to acquire (parts of) that information. Second-generation systems additionally exploit document relations, supporting users to reach a specific site, document, or the most authoritative one among many alternatives (i.e., navigational tasks). Third-generation systems blend results from different possibly multimodal vertical systems into a single SERP to support a user in performing transactional tasks.
We argue that generative retrieval systems can be seen as a new fourth generation of web search systems. Their synthesis of a single result “document” that condenses information from different sources relevant to some information need promises to reduce the users’ cognitive load compared to prior system generations that required users to condense the information themselves.444Sakai et al. (2011) had proposed to present lists of short automatically identified relevant information nuggets instead of complete documents in (2011), but they had not considered the aspect of condensing the nuggets to a single result. Additionally, the “synthesizing” nature of generative retrieval systems can conceivably be exploited to generate new pieces of information not contained in the retrieved sources, rendering the generative model itself another new source of information.
@@
While many of the search tasks addressed by generative retrieval systems may seem to be informational in nature, we still suggest to also separate the search tasks in a new category of synthetic search tasks. Complex needs like argumentative questions (Should society invest in renewable energy?) or decision-making questions (Should I get life insurance?) are simply not represented that well in Broder’s original categories. In contrast to informational tasks, the required information is hardly contained in some single document but rather spread across multiple documents; in contrast to navigational tasks, no single page is anticipated by the user to be reached; and in contrast to transactional tasks, the information condensation should be performed on the retrieval system side but not on the user side. Interestingly, as if already foreseeing generative retrieval, Broder even explicitly constrained informational queries and first-generation systems to static content: “The purpose of such [informational] queries is to find information assumed to be available on the Web in a static form. No further interaction is predicted, except reading. By static form we mean that the target document is not created in response to the user query.” (Broder, 2002, page 5).
The new fourth generation of web search systems supports synthetic search tasks and enables users to access a single, comprehensive generated answer document that can cover in-depth analyses of multiple perspectives on some complex information need. Although the Web may actually also offer some set of documents to satisfy complex needs, an ideal generative retrieval system can directly dynamically address them by retrieving relevant documents, synthesizing missing information, and condensing a coherent answer grounded in the retrieved sources.
@@ 2.4. A Taxonomy of Generative Retrieval
‘Generative retrieval’ or ‘generative IR’ are umbrella terms for a diversity of approaches that use generative models to solve retrieval tasks.555See also the recent SIGIR workshop on generative IR (Bénédict et al., 2023). Following Arora et al. (2023), Figure 3 categorizes these approaches into generation-augmented retrieval (GAR) and retrieval-augmented generation (RAG). Notably, GAR approaches create traditional list SERPs, while RAG approaches generate text SERPs.
In GAR approaches, generative models are used to enhance the traditional search architecture at indexing time or at query time. At indexing time, generative models can be used for augmenting documents (Nogueira et al., 2019; Gospodinov et al., 2023; MacAvaney et al., 2020; Formal et al., 2021; Zhuang and Zuccon, 2021) with confabulated or hallucinated content, or for replacing the standard indexing process with what are commonly termed ‘differentiable indices’ by, for instance, generating document identifiers like page titles (De Cao et al., 2021; Thorne, 2022; Chen et al., 2022), URLs (Ziems et al., 2023), or (structured) string identifiers (Zhou et al., 2023; Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022). At query time, generative models can be used for augmenting queries (MacAvaney et al., 2021; Alaofi et al., 2023), or for modeling relevance by, for instance, generating parts of existing documents from the query and retrieving the documents by string matching (Bevilacqua et al., 2022), by predicting a (re-)ranking directly (Sun et al., 2023), or by using special tokens as relevance signal (Nogueira et al., 2020; Ma et al., 2023; Qin et al., 2023; Zhuang et al., 2023).
In RAG approaches—the focus of our paper—, generative models are augmented with retrieval capabilities; either internally as ‘attention-level RAG’, where the context attended to during generation is retrieved concurrently (Guu et al., 2020; Jiang et al., 2022; Borgeaud et al., 2022), or externally as ‘prompt-level RAG’, where the retrieved context is inserted into the prompt. Orthogonally, one can distinguish the RAG variants retrieval-then-generation and generation-then-retrieval (cf. Section 2.2). Beyond GAR and RAG, generative models can also be used to directly generate a response without relying on retrieved information (Robinson and Wingate, 2023), i.e., as infinite indexes (Deckers et al., 2023). This may involve generating multiple candidates and selecting the best one or regenerating a new response conditioned on the previous ones (Yu et al., 2023). Moreover, an answer to a generative ad hoc request can also be the first turn of a conversational search (Salton, 1969; Radlinski and Craswell, 2017; Culpepper et al., 2018), where generative models have led to new tools (Zhang et al., 2021a; Miller et al., 2017) and dialog options (Zamani et al., 2020).
@@ 3. A User Model for Generative IR
The general structure of an information search process (Vakkari, 2016) as seen from the users’ perspective is shown in Figure 4 (top row). After formulating an information need and selecting and interacting with some search results, in a final synthesis step the users try to reach a satisfying outcome. While traditional list SERPs mainly assist the users during selection and interaction, the text SERPs of generative systems also directly encompass the synthesis step. Evaluating retrieval systems with respect to the information search process often relies on some model of user expectations and behavior. Yet, most current user models focus on list SERPs but not text SERPs. Thus, after preliminary considerations (Section 3.1), we explore how the information search process relates to generative approaches (Section 3.2). Afterwards, we follow the evaluation methodology proposed by Agosti et al. (2014): We define generative IR-oriented evaluation objectives for the search process (Section 3.3; shown in the bottom row of Figure 4) and we devise a user model corresponding to these objectives (Section 3.4). In Section 4, we then operationalize the user model.
The figure consist of two aligned rows. The upper row illustrates the information search process from a searcher’s perspective, with the four components of ‘formulation’, ‘selection’, ‘interaction’, and ‘synthesis’. The lower row illustrates evaluation objectives for generative IR systems, with the four components ‘prompting’, ‘retrieval’, ‘grounding’, and ‘presentation’. Three pairs of components are related: selection and retrieval, interaction and grounding, as well as synthesis and presentation. These three pairs represent the steps supported by generative IR systems.